a body travelling with uniform acceleration crosses two points A and B eith velocities 20 m/s and 30m/s respectivly, the speed of the body at mid point of A and B

Grade:12th pass

3 Answers

subhabrata ghosh
7 years ago
Let the acceleration of the car is = a   And the distance between A and B is = d
v-u2=2ad
v=30m/s and u=20m/s
2ad=302-202
ad=(900-400)/2=250
when the car is at the mid point of AB then let the speed of the car is v1
v12-202=2a(d/2)
v12=ad+400=250+400=650
therefore v1 is = 25.4950m/s
:) :)
Soumith Chowdary
7 years ago
Use v^2-u^2=2as where v=30 and u=20 and S is distance between A and B . Substitute S/2 in place of S In equation
SUVAN SETHY
3 years ago
Let the Distance between Ato B is  d 
and the mid point AB be C 
=> Vb ^2 = Va ^2+ 2ad => a = 250/d 
Similarly, 
Vc^2 = Va ^2 + 2a(d/2)
=> Vc= √650 = 25.49 m/s = 25.5 m/s (approx)